OpenStudy (anonymous):

Help me for a second time to answer my problem Elimination of Constants y=Cx+C^2+1 y=1+C1e^x+C2e^-2x y=9x^2 +Bx+c

OpenStudy (anonymous):

@ganeshie8 Can you please help me for a second time?

ganeshie8 (ganeshie8):

check the first equation once.... looks there is a typo ?

OpenStudy (anonymous):

ganeshie8 there is no typo that is different problems :)

OpenStudy (anonymous):

@dan815 Help me again? :)

OpenStudy (dan815):

go about it the same way as before find y': and y'': Then try to find a linear combination where the constants dissappear

ganeshie8 (ganeshie8):

y=Cx+C^2+1 y' = ?

OpenStudy (anonymous):

You're right I think my professor has a typo.

ganeshie8 (ganeshie8):

keep in mind, C is just a constant

OpenStudy (dan815):

are you sure its not cx^2 vanessa?

OpenStudy (anonymous):

dan815 yes i'm sure

OpenStudy (dan815):

ok continue

OpenStudy (anonymous):

I think the x in Cx is a sub

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

it helps if you can take a screenshot of equations and attach...

OpenStudy (anonymous):

@ganeshie8 Here's the photo

ganeshie8 (ganeshie8):

that doesn't help much lol both look same !

ganeshie8 (ganeshie8):

lets skip this and work the remaining problems first... ?

OpenStudy (anonymous):

What problems?

ganeshie8 (ganeshie8):

remaining two problems : y=1+C1e^x+C2e^-2x y=9x^2 +Bx+c

ganeshie8 (ganeshie8):

start by creating the table : y = 1+C1e^x+C2e^-2x y' = ? y'' = ?

ganeshie8 (ganeshie8):

see if can you find the first and second derivatives ^^

OpenStudy (anonymous):

Ahm ok :)

OpenStudy (anonymous):

HAHAHA I can derived that XD Lol

ganeshie8 (ganeshie8):

good, find the derivatives :)

OpenStudy (anonymous):

*I can't XD

OpenStudy (anonymous):

B'coz C is a constant XD

ganeshie8 (ganeshie8):

lol that makes our life easy

ganeshie8 (ganeshie8):

you can treat C as a number while differentiating

OpenStudy (anonymous):

hahaha You're right XD

ganeshie8 (ganeshie8):

give it a try

ganeshie8 (ganeshie8):

y = 1+C1e^x+C2e^-2x y' = (1+C1e^x+C2e^-2x)'

ganeshie8 (ganeshie8):

whats the derivative of 1 ?

OpenStudy (anonymous):

it's become 0 XD

ganeshie8 (ganeshie8):

see you do know how to differentiate it :) finish it off... !

ganeshie8 (ganeshie8):

y = 1+C1e^x+C2e^-2x y' = (1+C1e^x+C2e^-2x)' = 0 + ?

OpenStudy (anonymous):

Lol I just can diffirentiate only 1 XD

OpenStudy (anonymous):

the game is over XD

ganeshie8 (ganeshie8):

game just started :P

OpenStudy (anonymous):

wahahaha So Solve it ! XD Don't pressure me XD

ganeshie8 (ganeshie8):

derivative of e^x is also easy : \(\large (e^x)' = e^x\)

ganeshie8 (ganeshie8):

e^x is the only function whose derivative equals the function itself.

OpenStudy (anonymous):

Then? What's next ? :D

ganeshie8 (ganeshie8):

lol this is your game - you will need to sweat :P

ganeshie8 (ganeshie8):

Also \(\large (e^{ax})' = ae^{ax} \)

OpenStudy (anonymous):

Wahaah Lol

ganeshie8 (ganeshie8):

use them to find y' : y = 1+C1e^x+C2e^-2x y' = (1+C1e^x+C2e^-2x)' = 0 + ?

OpenStudy (anonymous):

Can we subtract the 1 and 2 eq. ? XD

OpenStudy (anonymous):

(y)2 ? y=2C1e^x+2C2e^-2x Y'=C1e^x+(-2C2e^-2x)

OpenStudy (anonymous):

y'-y=C1e^x

OpenStudy (anonymous):

Am I right? @ganeshie8

ganeshie8 (ganeshie8):

y' is right !

ganeshie8 (ganeshie8):

lets not rush to the final answer yet, cuz we have two constants c1 and c2.. that tells us that we need to differentiate it twice

OpenStudy (anonymous):

Ahh ok ! :)

ganeshie8 (ganeshie8):

y = 1+C1e^x+C2e^-2x y' = 0 + C1e^x -2C2e^-2x = C1e^x - 2C2e^-2x y'' = ?

ganeshie8 (ganeshie8):

differentiate y' again ^

ganeshie8 (ganeshie8):

y' = C1e^x - 2C2e^-2x y'' = ?